OpenStudy (anonymous):

use the four step process fo find f1(x) and then find f1(1), f1(2) and f1(3)....... f(x)= 8x/4+x

OpenStudy (anonymous):

f1(x) =8x/4+x

OpenStudy (anonymous):

all you have to do is substitute 1 for f1(1)

OpenStudy (anonymous):

f1(1)=8(1)/4+(1)=8/5

OpenStudy (anonymous):

can you do the rest?

OpenStudy (anonymous):

I can! thank you!

OpenStudy (anonymous):

your welcome, all the best, let me know if you have a problem...

OpenStudy (anonymous):

This is talking about f(x+h) to find f1(x)? It makes the equation different